C. Find the mean (𝝁𝑿), variance (πˆπ‘Ώ 𝟐) and standard deviation (πˆπ‘Ώ) of the sampling distribution of the means given the sample size, and the means and standard deviation of the population. Sample size n were randomly selected from the population



7. 𝑛 = 9

πœ‡ = 5.3

𝜎 = 3

Mean:


Variance:


Standard Deviation:


8. 𝑛 = 11

πœ‡ = 4.23

𝜎 = 5

Mean:


Variance:


Standard Deviation:


1
Expert's answer
2022-04-04T16:33:58-0400

We'll use the properties of sampling distributions of sample means.


7. Mean:

"\mu_{\bar x} =\mu=5.3."

Variance:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{3^2}{9}=1."

Standard deviation:

"\sigma_{\bar x}=\cfrac{\sigma}{\sqrt n}=\cfrac{3}{\sqrt 9}=1."


8. Mean:

"\mu_{\bar x} =\mu=4.23."

Variance:

"\sigma^2_{\bar x}=\cfrac{\sigma^2}{n}=\cfrac{5^2}{11}=2.27."

Standard deviation:

"\sigma_{\bar x}=\cfrac{\sigma}{\sqrt n}=\cfrac{5}{\sqrt {11}}=1.51."